# Qubit encoding for a mixture of localized functions

### v2

@@ Appendix E Extension to multidimensional cases
Here we establish a straightforward extension of the encoding scheme for one dimension to multidimensional cases. Although the extension is possible for arbitrary dimension, we focus on three-dimensional cases for practical interest.
@@ E.1 For generic basis functions
We introduce $x,y,$ and $z$ registers each of which is built up of $n_{q}$ qubits, forming the $3n_{q}$-qubit data register. We discretize each of the three directions for a cube into $N\equiv 2^{n_{q}}$ grid points similarly to the one-dimensional case. We assume the spacing $\Delta x$ of the grid points in each direction to be common. We want to encode a function in the three-dimensional space by using displaced localized function. To this end, by adopting $n_{f}$ localized functions for each direction, we build a basis set from product basis functions
@@
where $b_{\ell\mu}=0,\dots,n_{f}-1\ (\mu=x,y,z)$ specifies the localized function in $\mu$ space for the $\ell$th basis function in the three-dimensional space. The number of basis functions is thus $n_{\mathrm{loc}}=n_{f}^{3}.$ $a_{\ell\mu}$ represents the parameters in the localized function for the $\mu$ space collectively. We express the desired state as an LC of these functions as
@@ E.2 Probabilistic encoding using LFs as basis functions
The circuit for probabilistic encoding of $|\psi_{\mathrm{lc}}\rangle$ by using the product bases of discrete LFs is now constructed straightforwardly, as shown in Fig. 11. This circuit is a three-dimensional extension of $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}$ in Fig. 4(b). Although the nominal number of the constituent qubits in the data register is $3n_{q}$ (instead of $n_{q}$ in the main text) in this case, the discussion on the scaling of computational cost in the main text will suffer from no change.
@@ G.1 Relation between the Slater and Lorentzian function states
$|S;a,0\rangle_{n_{q}}$ $\xmapsto{\mathrm{QFT}}C_{S}(n_{q},a)\sum_{j=0}^{N/2-1}\frac{1}{\sqrt{N}}\sum_{k=0}^{N-1}\left(e^{-aj}\exp\frac{i2\pi jk}{N}+e^{-a(j+1)}\exp\frac{i2\pi(N-1-j)k}{N}\right)|k\rangle_{n_{q}}$
@@ G.2 Overlap between Lorentzian function states
Let us calculate the overlap $V(a,a^{\prime},k_{\mathrm{c}})\equiv\langle L;a,k_{\mathrm{c}}|_{n_{q}}|L;a^{\prime},0\rangle_{n_{q}}$ between two LFs. Since $|L;a,k_{\mathrm{c}}\rangle_{n_{q}}=T(k_{\mathrm{c}})|L;a,0\rangle_{n_{q}}=\mathrm{QFT}\cdot U_{\mathrm{shift}}(k_{\mathrm{c}})|S;a,0\rangle_{n_{q}},$ we obtain

### v3

@@ Appendix E Extension to multidimensional cases
Here we establish a straightforward extension of the encoding scheme for one dimension to multidimensional cases. Although the extension is possible for arbitrary dimension, we focus on three-dimensional (3D) cases for practical interest.
@@ E.1 For generic basis functions
We introduce $x,y,$ and $z$ registers each of which is built up of $n_{q}$ qubits, forming the $3n_{q}$-qubit data register. We discretize each of the three directions for a cube into $N\equiv 2^{n_{q}}$ grid points similarly to the one-dimensional case. We assume the spacing $\Delta x$ of the grid points in each direction to be common. We want to encode a function in the 3D space by using displaced localized function. To this end, by adopting $n_{f}$ localized functions for each direction, we build a basis set from product basis functions
@@
where $b_{\ell\mu}=0,\dots,n_{f}-1\ (\mu=x,y,z)$ specifies the localized function in $\mu$ space for the $\ell$th basis function in the 3D space. The number of basis functions is thus $n_{\mathrm{loc}}=n_{f}^{3}.$ $a_{\ell\mu}$ represents the parameters in the localized function for the $\mu$ space collectively. We express the desired state as an LC of these functions as
@@ E.2 Probabilistic encoding using LFs as basis functions
The circuit for probabilistic encoding of $|\psi_{\mathrm{lc}}\rangle$ by using the product bases of discrete LFs is now constructed straightforwardly, as shown in Fig. 12. This circuit is a 3D extension of $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}$ in Fig. 4(b). Although the nominal number of the constituent qubits in the data register is $3n_{q}$ (instead of $n_{q}$ in the main text) in this case, the discussion on the scaling of computational cost in the main text will suffer from no change.
@@ G.1 Relation between the Slater and Lorentzian function states
$|S;a,0\rangle_{n_{q}}$ $\xmapsto{\mathcal{F}_{\mathrm{quant}}}C_{S}(n_{q},a)\sum_{j=0}^{N/2-1}\frac{1}{\sqrt{N}}\sum_{k=0}^{N-1}\left(e^{-aj}\exp\frac{i2\pi jk}{N}+e^{-a(j+1)}\exp\frac{i2\pi(N-1-j)k}{N}\right)|k\rangle_{n_{q}}$
@@ G.2 Overlap between Lorentzian function states
Let us calculate the overlap $V(a,a^{\prime},k_{\mathrm{c}})\equiv\langle L;a,k_{\mathrm{c}}|_{n_{q}}|L;a^{\prime},0\rangle_{n_{q}}$ between two LFs. Since $|L;a,k_{\mathrm{c}}\rangle_{n_{q}}=T(k_{\mathrm{c}})|L;a,0\rangle_{n_{q}}=\mathcal{F}_{\mathrm{quant}}\cdot U_{\mathrm{shift}}(k_{\mathrm{c}})|S;a,0\rangle_{n_{q}},$ we obtain